OpenStudy (anonymous):

Simplify: cos 3x cos (2x+x) cos (2x+x) = cos 2x cos x - sin 2x sin x (cos^2(x) - sin^2(x))cos x - 2sin^2(x) cos x (cos^3(x) - sin^2(x) cos x) - 2sin^2(x) cos x Where do I go from here to reach one of these: A)cos x - 4 cos x sin^2(x) B) -sin^3(x) + 2 sin x cos x C) -sin^2(x) + 2 sin x cos x D) 2 sin^2(x) cos x - 2 sin x cos x

OpenStudy (anonymous):

there are other "double angle" formulas you might try for \(\cos(2x)\) that are less cumbersome

OpenStudy (anonymous):

like either \(1-2\sin^2(x)\) or \(2\cos^2(x)-1\)

OpenStudy (anonymous):

Such as: 2cos^2(x) - 1 1 - 2sin^2(x) So I just plug them in and hope it was the correct direction?

OpenStudy (anonymous):

Then again seeing that the answers have high powers of sin, then it'd be a good idea to use 1 - 2sin^2(x)?

OpenStudy (anonymous):

yeah that is what i would do

OpenStudy (anonymous):

cos 3x cos (2x+x) cos (2x+x) = cos 2x cos x - sin 2x sin x (1-2sin^2(x))cos x - 2sin^2(x) cos x cos x - 2sin^2(x)cos x - 2sin^2(x) cos x cos x - 4 cos x sin^2 x Like so?